Find the area of the region enclosed enclosed between the graphs of y=e^x and y=e^-x and the line x=ln2.
1
Expert's answer
2020-05-25T20:35:37-0400

"Area=\int_{ln(2)}^1(-ln(y)-ln(y))dy\\
\ \qquad \ = -2\int_{ln(2)}^1ln(y)dy\\
\ \qquad \ = -2y (ln(y)-1)|_{ln(2)}^1\\
\ \qquad \ = -2(ln(1)-1)+2ln(2)(ln(ln(2))-1)\\
\ \qquad \ = 2+2ln(2)(ln(ln(2))-1)\\"